I am trying to solve this problem and would appreciate some help:
An apparel shop sells skirts for $45 and blouses for $35 The entire stock is worth $51, 750.. Half the skirts and two thirds of the blouses are sold totaling $30,600. How many skirts and blouses are left in the store.

I have tried to formulate the equations. The text example uses multiplication to get the y correspondent to zero out when the two equations are added, but I can't get the right combo..............the answers are 260 skirts, 270 blouses

this is part of my attempts

45x+35y=51,750
.50(45x)+.66(35y)=30,600

22.5x+23.33y=30,600


Answer by richwmiller(17219)

Your first equation is ok.
I would use 2/3 instead of .66
45s+35b=51750
45s/2+2/3*35b=30600
We will want what is left not what was sold
b = 810, s = 520
This is the entire stock
We take 1/3 of the blouses and 1/2 of the skirts
810/3=B and 520/2=S
B=270 and S=260